OpenStudy (anonymous):

Suppose A = {3n | n E Z} and f: R --> R is defined by f(x) = X^(x). f is the characteristic function of set A. (a) What is f ({2n | n E Z})? (b) What is f^-1 ({2n | n E Z})?

OpenStudy (anonymous):

i am kind of lost A are integers divisible by 3, and \(f(x)=x^x\) right?

OpenStudy (anonymous):

see the screenshot. it is a confusing question

OpenStudy (anonymous):

oooh i see

OpenStudy (anonymous):

The characteristic function of the set A, \(\chi_A(x)\) is defined as: $$\chi_A(x)=1\quad\text{if } x\in A$$and: $$\chi_A(x)=0\quad\text{if } x\not{\in} A$$

OpenStudy (anonymous):

characteristic function means \(f(n)=1\) if \(n\in A\) and \(f(n)=0\) if \(n\notin A\)

OpenStudy (anonymous):

what @nerdguy2535 said

OpenStudy (anonymous):

The question is asking what is the image and pre-image of the set $$\{2n\mid n\in\mathbb{Z}\}$$ under the characteristic function.

OpenStudy (anonymous):

So for a), they want to know what the image of \(\{2n\mid n\in \mathbb{Z}\}\) is under f. In other words, if \(x\in \{2n\mid n\in \mathbb{Z}\}\), what are the possible answers for \(f(x)\)?

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

umm not yet...

OpenStudy (anonymous):

image of 2n would be {2, 4, 6,...} and preimage would be 2?

OpenStudy (anonymous):

i think the image is 1 and 0, since it is the characteristic function on integers divisible by 3

OpenStudy (anonymous):

1 and 0?

OpenStudy (anonymous):

the numbers that would get sent to 1 in 2n are the ones that are divisible by 6

OpenStudy (anonymous):

since they have to be divisible by 3 to get sent to 1, and all are even

OpenStudy (anonymous):

I am not sure if I understand it. can we go back to the beginning?

OpenStudy (anonymous):

sure the notation used is a little weird

OpenStudy (anonymous):

Suppose A = {3n | n E Z} and f: R --> R is defined by f(x) = X^(x). f is the characteristic function of set A. (a) What is f ({2n | n E Z})? // Do I need to find preimage and image of this?

OpenStudy (anonymous):

the set A is \(\{...-6,-3,0,3,6,9,...\}\)

OpenStudy (anonymous):

the function sends those numbers to 1 so \(f(-6)=1,f(-3)=1,f(0)=1,f(3)=1\) etc

OpenStudy (anonymous):

it sends everything else to \(0\) so \(f(1)=0,f(2)=0.f(\pi)=0\) etc

OpenStudy (anonymous):

I am sorry I am a little lost. where did f(-6) = 1 comes from?

OpenStudy (anonymous):

\[A=\{3n|n\in \mathbb{Z}\}\] so if \(n=-2\) then \(3n=-6\)

OpenStudy (anonymous):

A is the set of integers divisible by 3

OpenStudy (anonymous):

and since \(-6\) is divisible by 3, it is in A and so \(f(-6)=1\)

OpenStudy (anonymous):

ok what does it have to do with this "What is f ({2n | n E Z})?"

OpenStudy (anonymous):

the way i interpret the question, is what does \(f\) do to even numbers ?

OpenStudy (anonymous):

square it?

OpenStudy (anonymous):

f: R --> R is defined by f(x) = X^(x) <-?

OpenStudy (anonymous):

but then maybe i don't really understand the question i would say it sends numbers divisible by 6 to 1, all others to zero

OpenStudy (anonymous):

oh ok. I am sorry I still don't understand but thanks for trying to help

OpenStudy (anonymous):

yw, sorry i could not do better